# Gen2Seg: Generative Models Enable Generalizable Instance Segmentation

### v2

@@ Abstract
By pretraining to synthesize coherent images from perturbed inputs, generative models inherently learn to understand object boundaries and scene compositions. How can we repurpose these generative representations for general-purpose perceptual organization? We finetune Stable Diffusion and MAE (encoder+decoder) for category-agnostic instance segmentation using our instance coloring loss exclusively on a narrow set of object types (indoor furnishings and cars). Surprisingly, our models exhibit strong zero-shot generalization, accurately segmenting objects of types and styles unseen in finetuning. This holds even for MAE, which is pretrained on unlabeled ImageNet-1K only. When evaluated on unseen object types and styles, our best-performing models closely approach the heavily supervised SAM, and outperform it when segmenting fine structures and ambiguous boundaries. In contrast, existing promptable segmentation architectures or discriminatively pretrained models fail to generalize. This suggests that generative models learn an inherent grouping mechanism that transfers across categories and domains, even without internet-scale pretraining. Code, pretrained models, and demos are available on our website.
@@ 1 Introduction
$\text{COCO}_{\text{exc}}$
DRAM
EgoHOS
iShape
PIDRay
@@ 2.1 Generative Models for Perception
Figure 3: Our models assign similar colors to compositionally related parts of a scene. Vader’s mask and body (top), or the bowties and shirts (bottom) are separated by subtly different hues, while distinct colors partition unrelated parts such as his leg and the poles (top), or the dogs and text (bottom). This emerges without any part-level supervision, suggesting generative models learn hierarchical scene representations. More samples are provided in Figures 13 to 19.
@@ 3.1 Instance Segmentation as Image to Image Translation
Recent instance segmentation models predict sets of binary masks, each representing an object instance (Carion et al., 2020; Cheng et al., 2022). However, it is not obvious how to enable generative models, designed to map from $\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times 3}$, to easily decode to this style. Drawing inspiration from work on image-to-image translation (Isola et al., 2017; Zhu et al., 2017), we encode our ground truth as an RGB image with a unique color for each instance and black color for the background. We find that both the Stable Diffusion VAE and MAE models can decode these masks with effectively no loss in quality. Thus, we finetune our model by forwarding our RGB image (or its latent) without adding noise or masking, decoding the output, and optimizing in pixel space with respect to the RGB ground truth.
Unfortunately, one cannot simply assign each ground truth mask to a color and train the model to regress it since there are many possible accurate color assignments. Thus, we propose our instance coloring loss based on two key properties of an RGB segmentation mask. First, the color of all pixels within a mask should have low variance. Second, the color of a mask should not be predicted anywhere outside the mask. By emphasizing these two properties, we can learn a model of instance segmentation without needing to enforce specific colors for object masks. Simply, our finetuned model should ensure each instance is assigned a unique color that is consistent across its pixels. Our finetuning strategy is partially inspired by prior works (De Brabandere et al., 2017; Kong & Fowlkes, 2018) and thus we claim no novelty on it. We designed it to be simple, intuitive, and agnostic to model architecture with fast inference time.
@@ 4.3 Zero-Shot Promptable Instance Segmentation
SAM DINO-B MAE-B MAE-H SD MAE-H (COCO) SD (COCO) MAE-H (ClevrTex)
79.0 33.2 53.7 80.5 93.4 75.9 89.7 86.4
SD (ClevrTex) MAE-H (10 classes) SD (10 classes) MAE-H(5 Classes) SD (5 Classes) Sobel
88.5 86.9 92.5 83.1 91.7 65.0
Table 3: We evaluate the zero-shot edge AP for recall less than 20% of the edges synthesized from our models and baseline segmentations on BSDS500. Nearly all generative models outperform SAM, even when trained on polygonal mask edges from COCO. This suggests that our models’ ability to produce detailed groupings is due to the generative prior, and not any specific dataset.
@@ 4.4 Zero-Shot Edge Detection
We evaluate our segmentation features on the task of edge detection with the BSDS500 dataset (Martin et al., 2001). It is important to note that we are not simply trying to find all edges in the image, but only object boundaries. Following Kirillov et al. (2023), we simply apply a Sobel filter on the predicted features followed by nonmax suppression to thin the edges. We use the edge detection described above on the original image as a “weak” baseline and on the output of SAM’s AutoMaskGenerator as a “strong” baseline. We report AP for recall less than 20%. We explain this choice (and display the full precision-recall curves) in Appendix B.
@@ A.2 Evaluation Datasets
iShape: Many objects in the real world have fine, intricate structures. We evaluate our models’ ability to accurately segment these objects using the iShape dataset. iShape is an instance segmentation dataset which contains many occlusions and complex, fine structures such as wires or fences. We evaluate on the test set.
@@ Appendix B Additional Results and Experimental Details
Finally, in Table 7, we explore how deep we need to finetune our model to achieve strong results. We start with full finetuning and gradually reduce the layers trained to the last N layers in the model, plus the final projection to pixel space, freezing the rest of the model. We find that finetuning just the decoder leads to very similar performance to full finetuning. Additionally, we find that just one or two layers are enough to achieve reasonable performance. This suggests that the ”core” object grouping information is highly saturated in the decoder, especially the last layers. This makes sense because 1) since the rest of the model is frozen, only the representations that are available at the last N layers will actually have an effect on the output 2) these layers are the ones most responsible for synthesizing images, for which understanding object grouping is frozen.
@@
Decoder only (8 layers) 53.5 20.2 3.0
@@ D.2 $\text{COCO}_{\text{exc}}$
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
@@
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
@@
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
@@
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
@@
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM

### v3

@@ Abstract
By pretraining to synthesize coherent images from perturbed inputs, generative models inherently learn to understand object boundaries and scene compositions. How can we repurpose these generative representations for general-purpose perceptual organization? We finetune Stable Diffusion and MAE (encoder+decoder) for category-agnostic instance segmentation using our instance coloring loss exclusively on a narrow set of object types (indoor furnishings and cars). Surprisingly, our models exhibit strong zero-shot generalization, accurately segmenting objects of types and styles unseen in finetuning. This holds even for MAE, which is pretrained on unlabeled ImageNet-1K only. When evaluated on unseen object types and styles, our best-performing models closely approach the heavily supervised SAM, and outperform it when segmenting fine structures and ambiguous boundaries. In contrast, existing promptable segmentation architectures or discriminatively pretrained models fail to generalize. This suggests that generative models learn an inherent grouping mechanism that transfers across categories and domains, even without internet-scale pretraining. Please see our website for additional qualitative figures, code, and a demo.
@@ 1 Introduction
$\text{COCO}_{\text{exc}}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DRAM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EgoHOS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
iShape Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PIDRay Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 2.1 Generative Models for Perception
Figure 3: Our models assign similar colors to compositionally related parts of a scene. Vader’s mask and body (top), or the bowties and shirts (bottom) are separated by subtly different hues, while distinct colors partition unrelated parts such as his leg and the poles (top), or the dogs and text (bottom). This emerges without any part-level supervision, suggesting generative models learn hierarchical scene representations. See Figures 14 to 20 and Table 7 for more examples and results.
@@ 3.1 Instance Segmentation as Image to Image Translation
Recent instance segmentation models predict sets of binary masks, each representing an object instance (Carion et al., 2020; Cheng et al., 2022). However, it is not obvious how to enable generative models, designed to map from $\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times 3}$, to easily decode to this style. Drawing inspiration from work on image-to-image translation (Isola et al., 2017), we encode our ground truth as an RGB image with a unique color for each instance and black color for the background. We find that both the Stable Diffusion VAE and MAE models can decode these masks with effectively no loss in quality. Thus, we finetune our model by forwarding our RGB image (or its latent) without adding noise or masking, decoding the output, and optimizing in pixel space with respect to the RGB ground truth.
Unfortunately, one cannot simply assign each ground truth mask to a color and train the model to regress it since there are many possible accurate color assignments. Thus, we propose our instance coloring loss based on two key properties of an RGB segmentation mask. First, the color of all pixels within a mask should have low variance. Second, the color of a mask should not be predicted anywhere outside the mask. By emphasizing these two properties, we can learn a model of instance segmentation without needing to enforce specific colors for object masks. Simply, our finetuned model should ensure each instance is assigned a unique color that is consistent across its pixels. Our finetuning strategy draws on the same intuition as priors works that aim to cluster features for supervised instance segmentation (De Brabandere et al., 2017; Kong & Fowlkes, 2018). We designed it to enable a training strategy that is simple, intuitive, and agnostic to model architecture with fast inference.
@@ 4.3 Zero-Shot Promptable Instance Segmentation
COCO train image Annotated Masks SD (COCO)
Refer to caption Refer to caption Refer to caption
Model Edge AP Model Edge AP
DINO-B $33.2$ MAE-H (ClevrTex) $86.4$
MAE-B $53.7$ MAE-H (10 classes) $86.9$
Sobel $65.0$ SD (ClevrTex) $88.5$
MAE-H (COCO) $75.9$ SD (COCO) $89.7$
SAM $79.0$ SD (5 classes) $91.7$
MAE-H $80.5$ SD (10 classes) $92.5$
MAE-H (5 classes) $83.1$ SD $93.4$
Figure 5: We evaluate the zero-shot edge AP for recall less than 20% of the edges synthesized from our models and baseline segmentations on BSDS500. Nearly all generative models outperform SAM, even when trained on polygonal mask edges from COCO. Observe how in the figure above, even though the annotation has very noisy edges, SD (COCO) does not learn this and instead predicts smoother and more perceptually aligned boundaries. This suggests that our models’ ability to produce detailed groupings is due to the generative prior, and not any specific dataset.
@@ 4.4 Zero-Shot Edge Detection
We evaluate our segmentation features on the task of edge detection with the BSDS500 dataset (Martin et al., 2001). It is important to note that we are not simply trying to find all edges in the image, but only object boundaries. Following (Kirillov et al., 2023), we simply apply a Sobel filter on the predicted features followed by nonmax suppression to thin the edges. We use the edge detection described above on the original image as a “weak” baseline and on the output of SAM’s AutoMaskGenerator as a “strong” baseline. We report AP for recall less than 20%. We explain this choice (and display the full precision-recall curves) in Appendix B.
@@ A.2 Evaluation Datasets
iShape: Many objects in the real world have fine, intricate structures. We evaluate our models’ ability to accurately segment these objects using the iShape dataset. iShape is an instance segmentation dataset which contains many occlusions and complex, fine structures such as wires or fences. We evaluate on the test set using the PNG Cityscapes style annotations.
@@ Appendix B Additional Results and Experimental Details
8 layers (decoder only) 53.5 20.2 3.0
@@
Class No Compositionality MAE-H SD SAM
aeroplane (19) 17.5 10.3 19.7 48.6
bird (13) 15.6 14.2 20.3 36.3
bicycle (6) 19.5 28.4 27.6 46.6
bus (46) 12.7 7.5 9.5 56.0
cat (17) 10.7 22.5 29.0 39.1
cow (19) 11.4 11.0 17.3 26.9
dog (18) 11.0 19.0 25.9 34.2
horse (21) 9.2 10.7 17.8 26.2
motorbike (7) 15.5 21.1 29.1 54.0
person (24) 9.2 27.3 31.0 54.1
sheep (19) 14.4 10.0 15.5 27.1
train (52) 30.9 25.0 34.0 38.7
bottle (2) 53.4 28.0 34.4 53.1
car (30) 18.7 10.6 13.6 57.7
potted plant (2) 54.4 36.2 38.4 58.4
TV monitor (1) 55.5 51.3 57.4 66.9
Table 7: Emergent part-based grouping on Pascal-Part. We evaluate part-level grouping on the Pascal-Part dataset (part masks for PASCAL VOC 2012) by measuring IoU between discovered part groups and ground-truth part masks. “No Compositionality” is a baseline computed as the IoU between each part mask and its corresponding whole-object mask (e.g., cat head vs. whole cat), approximating performance with no part-level grouping. Bold indicates classes where MAE-H or SD improves by at least 5 IoU points over the baseline. Gray rows denote classes included in our Hypersim+VK2 training mix.
We also explore how deep we need to finetune our model to achieve strong results in Table 6. We start with full finetuning and gradually reduce the layers trained to the last N layers in the model, plus the final projection to pixel space, freezing the rest of the model. We find that finetuning just the decoder leads to very similar performance to full finetuning. Additionally, we find that just one or two layers are enough to achieve reasonable performance. This suggests that the ”core” object grouping information is highly saturated in the decoder, especially the last layers. This makes sense because 1) since the rest of the model is frozen, only the representations that are available at the last N layers will actually have an effect on the output 2) these layers are the ones most responsible for synthesizing images, for which understanding object grouping is frozen.
Finally, in Table 7, we quantitatively evaluate emergent part-based grouping on the Pascal-Part Chen et al. (2014) dataset, which annotates part-level masks for the PASCAL VOC 2012 set. Bolded values represent classes for MAE-H or SD where our model performed at least 5 points better than the baseline.
We compare against a baseline ”No Compositionality”, which is calculated by evaluating IoU between every part mask and its whole object mask. This represents the values close to what we would get if there was no part-level grouping at all (i.e. predicting the whole cat when ground-truth is just the cat head). We also report results from SAM to serve as a high-water mark and provide context for our results.
We see signs of part-level grouping for some classes such as bike, cat, cow, dog, horse, motorcycle, and person. This emerges without any supervision on part or whole-object masks of these categories, suggesting it is inherent to the generative model.
Classes colored gray are included in our Hypersim+VK2 mix; thus we do not expect to see any improvement on them as the whole-object supervision may overwrite knowledge of part-level groups.
@@ Appendix E Additional Qualitative Samples
Refer to caption Refer to caption DINO-B
MAE-H (5 classes) Refer to caption Refer to caption MAE-B
MAE-H (10 classes) Refer to caption Refer to caption MAE-H
MAE-H (COCO) Refer to caption Refer to caption MAE-H (ClevrTex)
SD (COCO) Refer to caption Refer to caption SD (ClevrTex)
SD (5 classes) Refer to caption Refer to caption SD (10 classes)
SD Refer to caption Refer to caption SAM
@@
Refer to caption Refer to caption DINO-B
MAE-H (5 classes) Refer to caption Refer to caption MAE-B
MAE-H (10 classes) Refer to caption Refer to caption MAE-H
MAE-H (COCO) Refer to caption Refer to caption MAE-H (ClevrTex)
SD (COCO) Refer to caption Refer to caption SD (ClevrTex)
SD (5 classes) Refer to caption Refer to caption SD (10 classes)
SD Refer to caption Refer to caption SAM
@@
Refer to caption Refer to caption DINO-B
MAE-H (5 classes) Refer to caption Refer to caption MAE-B
MAE-H (10 classes) Refer to caption Refer to caption MAE-H
MAE-H (COCO) Refer to caption Refer to caption MAE-H (ClevrTex)
SD (COCO) Refer to caption Refer to caption SD (ClevrTex)
SD (5 classes) Refer to caption Refer to caption SD (10 classes)
SD Refer to caption Refer to caption SAM
@@
Refer to caption Refer to caption DINO-B
MAE-H (5 classes) Refer to caption Refer to caption MAE-B
MAE-H (10 classes) Refer to caption Refer to caption MAE-H
MAE-H (COCO) Refer to caption Refer to caption MAE-H (ClevrTex)
SD (COCO) Refer to caption Refer to caption SD (ClevrTex)
SD (5 classes) Refer to caption Refer to caption SD (10 classes)
SD Refer to caption Refer to caption SAM
@@
Refer to caption Refer to caption DINO-B
MAE-H (5 classes) Refer to caption Refer to caption MAE-B
MAE-H (10 classes) Refer to caption Refer to caption MAE-H
MAE-H (COCO) Refer to caption Refer to caption MAE-H (ClevrTex)
SD (COCO) Refer to caption Refer to caption SD (ClevrTex)
SD (5 classes) Refer to caption Refer to caption SD (10 classes)
SD Refer to caption Refer to caption SAM